EtiQuestion 4: Merrimack Co. is a U.S. firm that is consideringthe location of a new investment project. With expected return 20
Answer #1

4 a ) 70% is invested in existing business and rest 30% in new project which can be either in UK or in US

As the return of a portfolio is the weighted return of the constituent projects and

The standard deviation of an n security/project portfolio is given by

Op = W;*W, *0; * 0;* Pij

where Wi is the weight of security/project i in the portfolio

Sigma(i) is the standard deviation of returns of security/project i and

rho (i,j) is the correlation coefficient between returns of security/project i and security/project j

If the project is located in UK

Return of the portfolio = 0.70 * 20% +0.30 *25% = 21.5%

Standard deviation =(0.7^2*0.1^2+0.3^2*0.11^2+2*0.7*0.3*0.1*0.11*0.02) ^(0.5)

=(0.006081)^0.5

=0.077983 = 7.7983%

If the project is located in USA

Return of the portfolio = 0.70 * 20% +0.30 *25% = 21.5%

Standard deviation =(0.7^2*0.1^2+0.3^2*0.09^2+2*0.7*0.3*0.1*0.09*0.8) ^(0.5)

=(0.008653)^0.5

=0.093022 = 9.3022%

b) As can be seen from the above risk and return situations , the returns of the overall business remains the same in case of whether the project is done in UK or USA. However, the risk of the portfolio reduces when the project is located in UK. This happens despite the higher risk of the standalone UK project as compared to the USA project ( standard deviation of UK project is 0.11 whereas for USA project , it is 0.09) .

The portolio risk reduces because of International diversification of risk reflected in the low correlation coefficient of returns of the US project and UK project. (0.02) . The correlation coefficient signifies that the returns of the existing project in USA and the new project in UK are almost uncorrelated. i.e. not affected by each other. As a result , the portfolio risk decreases